OpenStudy (anonymous):

Find an equation of the tangent line to the curve y=(e^x)(sinx) at the point (0,0).

OpenStudy (callisto):

1) find y' 2) find y' at (0,0). This is the slope 3) You have the slope, you have a point (0,0). You can form an equation by point-slope form

OpenStudy (anonymous):

ok y'=e^xsinx+e^xcosx

OpenStudy (anonymous):

when x=0 y=1

OpenStudy (anonymous):

and when y=0 x=some random number

OpenStudy (anonymous):

@Callisto

OpenStudy (callisto):

No.... You need to find the slope of tangent line at (0,0) So, you need to sub (0,0) into y' to get the slope.

OpenStudy (anonymous):

how do you plug a point into the equation. if you put in zero then that means you will get zero

OpenStudy (callisto):

Try.

OpenStudy (callisto):

Sorry, plug x=0 into y'

OpenStudy (anonymous):

ok i did and got 1

OpenStudy (callisto):

1 is the slope m. \[y-y_0 = m(x-x_0)\]\((x_0,y_0) = (0,0)\)

OpenStudy (phi):

fyi, here is a graph

OpenStudy (anonymous):

so the equation would be y=x

OpenStudy (callisto):

Yes.